### v1

# RAGAR: Retrieval Augment Personalized Image Generation Guided by Recommendation

Abstract
Personalized image generation is crucial for improving the user experience, as it renders reference images into preferred ones according to user visual preferences. Although effective, existing methods face two main issues. First, existing methods treat all items in the user historical sequence equally when extracting user preferences, overlooking the varying semantic similarities between historical items and the reference item. Disproportionately high weights for low-similarity items distort users’ visual preferences for the reference item. Second, existing methods heavily rely on consistency between generated and reference images to optimize the generation, which leads to underfitting user preferences and hinders personalization. To address these issues, we propose Retrieval Augment Personalized Image GenerAtion guided by Recommendation (RAGAR). Our approach uses a retrieval mechanism to assign different weights to historical items according to their similarities to the reference item, thereby extracting more refined users’ visual preferences for the reference item. Then we introduce a novel rank task based on the multi-modal ranking model to optimize the personalization of the generated images instead of forcing depend on consistency. Extensive experiments and human evaluations on three real-world datasets demonstrate that RAGAR achieves significant improvements in both personalization and semantic metrics compared to five baselines.
1 Introduction
Personalized image generation has been widely applied in scenarios such as advertising systems and chat software. It aims to render reference images into preferred ones based on user visual preferences. Existing image generation methods [???] often produce similar outputs for different users with similar input, failing to meet users’ need for personalization. As a personalized generation method, [?] extracts user preferences from user historical sequences by treating each item equally, ignoring the varying semantic similarities between the reference item and historical items, such as semantic differences in categories. This disproportionately amplifies the influence of irrelevant items, causing the extracted user preference to deviate from the true visual preferences on the reference item. Fig. 1(a) shows a comparative example, if a user historical sequence includes electronic products and clothes, and the reference item is ”short sleeves”, an overly simplistic approach might generate a striped T-shirt with the Apple logo—-an outcome that misses the user’s true preferences. By incorporating semantic-based retrieval, more relevant visual preferences, such as ”clothing, sport, pink”, can be captured. Hence, we propose the assumption that filtering the items semantically related to the reference item enhances user preference modeling.
Figure 1: A comparative example shows the effect of retrieval and recommendation by contrasting generated results with and without them.
Another challenge lies in defining and evaluating personalization quantitatively. Current methods rely on either human evaluation or large multi-modal models [??], both of which are resource-intensive. Recent work [?] evaluates and rectifies generated images heavily based on their pixel similarity to the reference image, resulting in limited personalization. Inspired by advances in recommendation systems, which excel in learning and assessing personalized preferences, we propose using multi-modal recommendation models to evaluate generated images, shown in Fig. 1(b). These models provide an efficient and effective way to measure personalization. Moreover, by combining models and metrics from the recommendation domain, we validate in Sec. 2 that: 1) items selected through retrieval are more closely aligned with user preferences than randomly selected items; 2) retrieving items semantically related to the reference item enhances the learning of user preferences compared to irrelevant items.
Hence, the paper proposed the Retrieval Augment Personalized Image GenerAtion guided by Recommendation (RAGAR), comprising three key modules: 1) Retrieval Module contains a correlation unit to calculate the semantic similarity between historical and reference items and a fusion unit to fuse the filtered visual features with dynamic weight. 2) Generation module applies a balance calibrator to combine general user preferences derived from the fine-tuned LLM and the retrieval-augmented user preferences to produce images. 3) Reflection module calculates the loss to balance semantics and personalization, guided by the discriminator.
In summary, the key contributions are as follows:
  • We are the first to emphasize the relationship between historical items and the reference item. The corresponding assumptions are supported by data analysis on real-world datasets in Sec. 2.
  • We propose a novel model RAGAR to generate personalized images. We utilize the retrieval assumption realized by calculating the semantic similarity between items to enhance semantic consistency. Moreover, we design a discriminator to refine and assess the generated images, ensuring higher personalization quality.
  • Experiments on three real-world datasets demonstrate that our proposed method can achieve significant improvements in both personalization and semantic alignment compared to five competing methods.
2 Assumptions Validation
In this section, we validate our assumptions through data analysis on three real-world datasets: POG, ML-latest, and Stickers. [^1] Dataset descriptions are provided in Sec. 5.1. Details can be found in Table 1. To evaluate the effectiveness of the retrieval method, we use the multi-modal recommendation model [?] as the ranking model. We compare three types of user historical sequences: 1) ‘Ret’ contains the top-$k$ items selected using the retrieval method described in Sec. 4.3. 2) ‘Exp-Ret’ (Expanded Retrieval) contains the items ranked between $(k,2k]$ by the retrieval method. 3) ‘Random’ contains $k$ randomly selected items. To assess how well these sequences capture user preferences, we use Recall@10 and NDCG@10 as evaluation metrics.
Figure 2: Assumption validation results across three datasets.
The results shown in Fig. 2. ‘Random’ has the lowest values in metrics, demonstrating that randomly selected items are less effective for preference modeling. Our first assumption holds because both ‘Ret’ and ‘Exp-Ret’ achieve higher metrics, suggesting that items related to the reference items are crucial for modeling user preferences. Moreover, both metrics of ‘Ret’ are higher than those of ‘Exp-Ret’, which implies that the second assumption holds. Therefore, both of our assumptions hold in practical applications, motivating us to retrieve higher ranked items to enhance the user preferences modeling.
3 Related Work
Figure 3: The overview framework of the proposed RAGAR model.
Personalized Image Generation.
Advances in diffusion models (DMs) [?] have enabled high-quality image generation [????]. However, personalized image generation where user preferences guide the generation process, remains underexplored. Methods like [?] and [?] rely on natural language instructions, while [???] combine user interactions and instructions to create personalized advertisements or product images. Recent works bridge DMs and large language models (LLMs) [???], such as [?]. Building on this foundation, [?] incorporates user historical sequences to extract personalized features. However, [?] relies solely on consistency loss for optimization, overfitting the reference image’s features while neglecting personalized preferences. End-to-end multi-modal large language models (MLLMs) [???] show promise in personalizing images from textual instructions often overlook fine-grained preferences. Moreover, widely used evaluation metrics like FID [?], SSIM [?], LPIPS [?], and CLIP score [?] do not align well with human judgment, making them inadequate for assessing personalization. Meanwhile, evaluating personalization manually or using MLLMs [??] is inherently resource-intensive.
Recommendation Systems.
The goal of recommendation systems is to deliver personalized content that meets user preferences. These systems often rely on retrieval [??] and recommendation [???] techniques. Under the multi-modal setting, approaches like [?] integrate textual and visual features for enhanced commonsense reasoning, while [?] extends [?] with image understanding adapters. Specifically, [?] captures user-item relationships across modalities, enabling precise preference modeling for personalized recommendations.
4 Method
In this section, we formulate our task, introduce key notation, and outlines the RAGAR mainstream.
4.1 Task Formulation
We introduce the notation and task formulation for personalized image generation. For the user $u$, let the historical sequence be denoted as $\mathbf{S}_{u}=\{I_{1},I_{2},\cdots,I_{N-1}\}$, and let $I_{N}$ represent the reference item. Personalized image generation aims to produce an image that aligns with the user’s personalized preferences and preserves semantic consistency with $I_{N}$. Under the multi-modal setting, each item $I_{j}$ consists of an image and a text description, denoted as $\{img_{j},txt_{j}\}$.
4.2 RAGAR
Our proposed method, RAGAR, is illustrated in Fig. 3. To reduce the impact of irrelevant items, we introduce the retrieval module that calculates semantic similarity scores between the reference item and historical items. Then, the module fuses the visual features of the items with the weight corresponding to the scores to produce retrieval-augmented preference features. Next, the generation module employs a fine-tuned LLM to derive general preference features from multi-modal historical sequences, corrected by the retrieval-augmented preference. A diffusion-based generator then generate images conditioned by corrected preference. Finally, we design a reflection module to evaluate the generated images, balancing personalization and semantics to ensure high-quality results.
4.3 Retrieval Module
The retrieval module aims to integrate visual features of historical items that are semantically relevant to the reference item, including the correlation unit and the fusion unit.
Correlation Unit.
The correlation unit transforms image information from historical items into textual features to compute semantic similarity scores with the reference item. Given the historical sequence $\mathbf{S}_{u}$ and reference item $I_{N}$ for the user $u$, we extract the textual description from the item images using a caption model (e.g., BLIP-2 [?]):
$cap_{i}=\textbf{Caption}(img_{i}),i=0,1,\cdots,N$ (1)
where $cap_{i}$ represents the caption of item $I_{i}$. Next, the textual description is processed into high-dimensional semantic features with a text encoder, such as CLIP [?]:
$\mathbf{E}^{sem}_{i}=\mathbf{Encoder}(cap_{i})$ (2)
where $\mathbf{E}^{sem}_{i}$ denotes the high-level semantic representation of each caption. Then, we calculate the semantic similarity between historical items and the reference item using cosine similarity:
$s_{i}=cosine(\mathbf{E}^{sem}_{i},\mathbf{E}^{sem}_{N}),i=0,1,\cdots,N-1$ (3)
where $s_{i}$ represents the similarity score for each historical item relative to the reference item. The similarity scores are sorted in descending order, and the top-$k$ items form the retrieval sequence $\mathbf{S}^{ret}_{u}$, capturing the most relevant items in the user’s history.
Fusion Unit.
The fusion unit integrates preference features from the retrieval sequence by computing a weighted sum of visual features based on their semantic similarity scores, emphasizing high-similarity items while filtering out low-similarity ones. For the retrieval sequence $\mathbf{S}^{ret}_{u}$, we extract the visual features using a visual encoder:
$\mathbf{E}^{vis}_{i}=\mathbf{Encoder_{V}}(img_{i}),i=0,1,\cdots,N-1$ (4)
where $\mathbf{E}^{vis}_{i}$ denotes the high-level visual representation of each image. Then, the retrieval-augmented preference feature is computed as a weighted sum of these visual features and their corresponding similarity scores:
$\mathbf{P}^{ret}=\sum_{i=1}^{k}\mathbf{E}^{vis}_{i}\cdot s^{\prime}_{i}$ (5)
where $\mathbf{P}^{ret}$ represents the retrieval-augmented preference feature for the user $u$, and $s_{i}^{\prime}$ denotes normalized similarity scores.
4.4 Generation Module
The generation module bridges general user preferences with retrieval-augmented preferences to produce personalized images. It operates in three stages: 1) extracting general preference features using an LLM, 2) correcting the general preference features with the balance calibrator, and 3) generating personalized images using a diffusion-based generator.
Prompt Construction.
To capture user preferences from interactions, we transform interacted items into structured textual descriptions suitable for LLM analysis. Specifically, we adopt the approach from [?] that summarizes items with concise keywords to enhance interpretability and reduce extraneous information. We construct a prompt $p_{k}$ (detailed in supplementary material) populated with item captions and text to extract concise keywords with an LLM, denoted as $\mathbf{LLM}_{k}$:
$w_{i}=\mathbf{LLM}_{k}(p_{k}(cap_{i},txt_{i})),i=0,1,\cdots,N-1$ (6)
where $w_{i}$ is the keyword list for item $I_{i}$. We filter out low-frequency keywords and retain the top-$n$ keywords as $w^{\prime}$. Next, we construct a prompt $p_{g}$ to capture the general user preference. To enhance the expressivity of the LLM, denoted as $\mathbf{LLM}_{g}$, for image generation, we expand its vocabulary by incorporating $L$ additional special tokens ${\left[IMG\{i\}\right]}_{i=1}^{L}$. The tokens and keywords $w^{\prime}$ are then populated to $p_{g}$ (detailed in supplementary material) to extract preferences.
General preference Extraction.
Given the prompt $p_{g}$, the output of $\mathbf{LLM}_{g}$ can be divided into two embeddings:
$[\mathbf{E}^{txt};\mathbf{E}^{img}]=\mathbf{LLM}_{g}(p_{g})$ (7)
where $\mathbf{E}^{txt}$ represents text-related embedding and $\mathbf{E}^{img}\in\mathbb{R}^{L\times d}$ encodes image-related features. To bridge the gap between textual description and visual representation in LLMs, we adopt the Modal Mapper to align $E^{img}$ with the image space following [?], a 4-layer encoder-decoder transformer with trainable queries and two linear layers. Then we concatenate the text-related embedding and the image-related embedding and obtain the multi-modal feature.
$\mathbf{E}^{m}=[\mathbf{E}^{txt};\mathbf{ModalMapper}(\mathbf{E}^{img})]$ (8)
To maximize keyword utilization, we use the text encoder to generate keyword feature $\mathbf{E}^{key}$. We then employ a cross-attention layer to integrate $\mathbf{E}^{m}$ with $\mathbf{E}^{key}$:
$\mathbf{P}^{gen}=[Attn(\mathbf{E}^{m},\mathbf{E}^{key},\mathbf{E}^{key});\mathbf{E}^{key}]$ (9)
where $Attn$ denotes the cross attention layer to fuse features. In this way, we obtain the general preference feature $P^{gen}$.
Balance Calibrator.
As mentioned in Sec. 4.3, the retrieval-augmented preference focuses primarily on item features associated with the reference item. To narrow the gap between the retrieval-augmented preference feature $\mathbf{P}^{ret}$ and the global preference feature $\mathbf{P}^{gen}$, we minimize the calibrator loss between them:
$\mathcal{L}_{\mathrm{cal}}=\|\mathbf{P}^{gen}-\mathbf{P}^{ret}\|_{2}^{2}$ (10)
This adjustment ensures that $P^{gen}$ reflects both the general and retrieval-augmented preferences.
Image Generation.
With the corrected preference, we utilize the diffusion-based generator to generate images. During training, we sample $r$ random noises $\{\epsilon_{t}\}_{t=1}^{r}\sim\mathcal{N}(\mu,\sigma^{2}I)$ and apply reparameterization trick to enable gradient backpropagation through the sampling process:
$\mathbf{z_{t}}=\mu+\sigma\odot\epsilon_{t}$ (11)
The generator then produces $r$ images using the policy gradient update strategy discussed in Sec. 4.5. Additionally, the keyword feature $\mathbf{E}^{key}$ is used to generate a reference-like image, which serves as input to the ranking model for comparison with the personalized images during the reward computation.
$\begin{split}v^{key}&=Unet(\mathbf{E}^{key})\\ v^{gen}_{t}&=Unet(\mathbf{P}^{gen}+\mathbf{z_{t}}),t=0,1,...,r\end{split}$ (12)
where $v^{key}$ represents the image generated using the keyword features, and $v^{gen}$ represents the image generated using the general preference features.
4.5 Reflection Module
As the diffusion-based generator, propagating gradients backwards from the generated images is challenging [?]. To address this issue, we design a two-part reflection module that balances personalization and semantics in the generated images, ensuring both user alignment and semantic consistency.
Personalization Reflection
To provide personalized feedback on generated images, we leverage a pre-trained multi-modal ranking model (RM) to evaluate images, reducing the reliance on expensive manual labeling. Given a set of images $\{v^{ref},v^{key},v^{gen}_{t}\}$, where $v^{ref}$ represents the reference image. By substituting visual features from this set into the sequence of original items, RM assigns the scores and rankings for each item:
$\{\rho,rk\}=\textbf{RM}(S_{u},v)$ (13)
where $\rho\in\{\rho^{ref},\rho^{key},\rho^{gen}_{t}\}$ is the rank score, and $rk\in\{1,2,3\}$ denotes the rank derived from the score. Intuitively, images that align better with the user’s personalization preference are expected to achieve higher scores and lower ranks. We then employ a reward function to guide the model toward generating images that reflect personalization preferences. Inspired by the policy gradient method [?], we accumulate rewards for each noise $\epsilon_{t}$:
$\mathcal{R}_{t}=\sum_{c\in\{ref,key\}}\max\left(0,\rho^{c}-\rho^{gen}_{t}\right)+\delta$ (14)
$\mathcal{L}_{rank}=-\frac{1}{r}\sum_{t=1}^{r}\log_{\pi_{\theta}}(\mathbf{P}^{gen}+\epsilon_{t})\cdot\mathcal{R}_{t}$ (15)
where $\delta$ is a margin that encourages the model to generate higher-quality images, $\log_{\pi_{\theta}}$ computes probability density of the generated embedding under the Gaussian distribution.
Semantic Reflection
To enhance the semantic consistency, we minimize the distance between the mapper’s output in the generation module, $\mathbf{E}^{m}$, and the semantic features of the reference image, $\mathbf{E}^{sem}_{N}$ using the semantic loss:
$\mathcal{L}_{\mathrm{sem}}=\|\mathbf{E}^{m}-\mathbf{E}^{sem}_{N}\|_{2}^{2}$ (16)
Combined with the calibrator loss, generated images align both semantically and visually with the reference.
Joint Reflection
We employ LoRA [?] to update a limited set of parameters in LLM, the Modal Mapper and the attention fusion components. The overall loss is defined as:
$\mathcal{L}=\alpha\cdot\mathcal{L}_{cal}+\beta\cdot\mathcal{L}_{\mathrm{rank}}+\gamma\cdot\mathcal{L}_{\mathrm{sem}}$ (17)
where $\alpha$, $\beta$ and $\gamma$ are adjustable weighting hyper-parameters.
Features POG ML-latest SER30K
#Users 1000 4689 2230
#Items 19242 9742 30739
Table 1: The characteristics of experimental datasets.
5 Experimentation
We evaluate RAGAR on three different scenarios: commodity, movie poster and sticker. Our experiments aim to answer the following research questions:
  • RQ 1: How does RAGAR perform compared to other generative methods in quantitative evaluation metrics?
  • RQ 2: Dose RAGAR outperform baseline methods in human evaluation?
  • RQ 3: What is the impact of each module of RAGAR on performance? Specifically, how do the retrieval module and the reflection module contribute in the generation?
  • RQ 4: Can personalized images generated by RAGAR improve the performance in recommendation systems?
5.1 Experimental Settings
Datasets.
We utilize three real-world datasets to generate personalized images across different scenarios: POG [^2] https://github.com/wenyuer/POG.,ML-latest [^3] https://grouplens.org/datasets/movielens.,SER30K [^4] https://github.com/nku-shengzheliu/SER30K.. POG is a multi-modal dataset of fashion clothing with user interaction history. Due to the large size, we randomly select a subset of users and interactions. ML-latest is a benchmark movie dataset with user ratings. We collect corresponding posters from IMDB [^5] https://www.imdb.com. and split user historical sequences to ensure each contains 20 interactions. SER30K is a large-scale sticker dataset where each sticker is categorized by theme and annotated with an associated emotion label. We merge two random themes into user historical sequences and randomly select a reference sticker from other themes. Dataset statistics are summarized in Table 1.
Datasets Methods Personalization Semantic Alignment
[…]
POG GLIDE -54.02 13.90 59.01 49.01 17.24 17.52 59.17 58.35 27.14
SD-v1.5 -25.88 15.02 62.97 55.46 12.53 24.18 63.72 63.50 12.39
TI -25.64 15.64 63.79 53.68 15.67 24.14 67.83 62.40 15.23
LaVIT -51.09 15.40 63.46 48.77 22.65 24.07 72.04 56.42 24.79
PMG -22.96 14.82 56.74 55.76 5.19 18.48 57.00 63.89 5.02
RAGAR -19.77 15.79 66.88 55.93 17.99 24.28 74.55 59.29 19.08
ML-latest GLIDE -3.20 13.27 29.52 60.34 17.99 18.21 42.31 61.88 26.22
SD-v1.5 -1.97 12.85 31.07 59.74 14.02 18.38 53.59 58.58 14.51
TI -2.58 14.42 32.85 59.33 14.29 25.25 54.17 59.35 14.79
LaVIT -1.54 14.80 34.78 57.01 18.16 21.04 57.56 56.35 19.64
PMG -0.11 15.07 41.80 53.83 7.34 14.30 43.85 53.51 7.29
RAGAR 0.01 16.20 43.45 56.56 15.14 19.04 53.24 56.30 15.75
Sticker GLIDE -0.86 12.27 49.52 59.78 18.00 16.47 48.97 67.88 28.23
SD-v1.5 -0.93 12.48 51.08 59.45 17.38 18.46 49.61 68.41 16.43
TI -0.83 13.24 50.85 59.48 16.76 17.26 50.14 67.17 15.72
LaVIT -0.75 12.64 53.05 57.82 24.41 21.12 58.99 62.27 24.43
PMG -0.73 13.31 50.97 59.93 4.94 18.72 50.92 68.81 4.73
RAGAR 0.99 14.97 55.25 56.71 25.61 17.69 64.41 61.38 27.48
Table 2: Quantitative performance comparison on three datasets in terms of personalization and semantic alignment. The best performance is bold while the second-best is underlined.
Comparison Methods.
We compare RAGAR with five generative baselines, including three DM-based models Glide [?], SD [?] and TI [?], and two LLM-based models LaVIT [?] and PMG [?].
Evaluation Metrics.
We compare RAGAR with baselines through the following quantitative metrics: To evaluate personalization, we calculate the CLIP Personalization Score (CPS) and the CLIP Personalization Image Score (CPIS), which measure the similarity between the generated images and the text description or images representing user preferences, respectively. In addition, we calculate the LPIPS [?] and SSIM [?] to quantify the perceptual similarity. Furthermore, we compute the rank change $\mathbf{\Delta R}$ between the original and generated images. This metric emphasizes improvements for higher-ranked items by diminishing the weight for lower-ranked ones. A larger $\mathbf{\Delta R}$ indicates greater personalization in the image.
$\mathbf{\Delta R}=\frac{rk_{ori}-rk_{gen}}{1+rk_{ori}}$
where $rk$ denoted the rank determined by the ranking model. To evaluate semantic alignment, we calculate CLIP Score (CS), CLIP Image Score (CIS), LPIPS [?] and SSIM [?] for reference images.
Parameter Settings.
To make a fair comparison, all baselines are tuned with a fixed learning rate of $1e^{-5}$ and stable diffusion 1.5 is used as the image generator. For RAGAR, we set the learning rate at $1e^{-5}$. The number of retrieval items is fixed to 5. The number of noise is set at 3. All experiments are conducted on a single NVIDIA-A100 GPU.
5.2 Performance Comparison (RQ1)
The comparison between RAGAR and baselines is shown in Tab. 2. We can draw the following conclusions. First, traditional DM-based methods perform poorly on personalization metrics. Specifically, Glide and SD are the worst performers across five personalization metrics in all datasets. This is because both methods rely on the CLIP model to generate images, which struggles to capture deeper associations between texts. TI introduces a word embedding approach based on SD to capture user preferences. The stylized word embedding significantly improves personalization performance across all datasets (e.g., for $\Delta R$: -25.64 vs. -54.02, -2.58 vs. -3.20, -0.83 vs. -0.86). In terms of semantic alignment, all three methods perform well, and Glide achieves optimal results on SSIM across all datasets. This indicates that traditional methods tend to focus on reconstructing reference images.
Second, LLM-based methods demonstrate better performance, benefiting from their ability to understand and analyze text comprehensively. LaVIT, a multi-modal vision-language transformer, is designed to align textual and visual features effectively. Although it outperforms DM-based methods, it still falls short in personalization ($\Delta R$: -51.09 in the POG dataset). PMG, an LLM-based personalized image generation method, leverages LLMs to extract user preferences. This approach significantly enhances personalization performance, achieving optimal results compared to DM-based baselines. However, it heavily relies on the consistency with reference images, limited the personalization performance.
Third, our proposed method, RAGAR, achieves state-of-the-art (SOTA) performance due to the retrieval for sequence and the rank model for training across the three datasets. The performance boost is more pronounced on the ML dataset compared to POG and Sticker, likely because the amount of data in the front is significantly larger than that in the back. Compared to PMG, the second-best model, our method enhances consistency by semantically retrieving sequences and introducing evaluation metrics that better align with human judgment, resulting in stronger personalization performance.
Methods POG ML-latest Sticker
Per. Sen. Per. Sen. Per. Sen.
ORI 2.26 / 2.16 / 2.18 /
PMG 2.08 1.62 2.1 1.94 2.24 1.88
RAGAR 1.66 1.46 1.74 1.66 1.58 1.12
Table 3: Human evaluation comparison on both personalization and semantic alignment.
5.3 Human Evaluation (RQ2)
To assess the personalization and semantics of RAGAR in real scenarios, we conduct human evaluation to compare it with PMG and original images. We invited 50 volunteers and set up two types of sorting tasks, with 50 cases for each dataset, covering personalization and semantics. For personalization, given 5 historical images along with the generated images, participants are asked to sort images in descending order of likelihood that the user would click them, based on historical click data. For semantics, given the reference image and generated images, participants are asked to sort them in descending order of how well they convey the same meaning as the reference image. As shown in Tab. 3, RAGAR outperforms the baselines on both metrics. This indicates RAGAR better reflects user preference while preserving the semantics.
Figure 4: Ablation study of the retrieval module and the rank rewards.
Figure 5: Examples of image generation across three datasets.
5.4 Ablation Study (RQ3)
We study the effectiveness of key components of RAGAR in the POG dataset, including retrieval and reflection module.
Effect of retrieval module.
We assess the impact of retrieval module on improving semantic consistency and capturing user preferences by excluding it during training. Fig. 4(a) shows that: 1) noisy items in the historical sequence diminish the model’s ability to capture user preferences and interfere with semantics; 2) the retrieval module selects items relevant to user preferences, thereby improving the performance. Furthermore, we investigate the effect of varying retrieval number $k$ during training. The results reveal that $k=5$ yields the best results when historical sequence length is 20. At $k=0$ (no retrieval), performance is poor, while increasing $k$ initially improves performance but declines beyond $k=5$ due to reintroduced noise and redundancy.
Effect of rank rewards.
We investigate the role of rank rewards by excluding it while retaining the other loss functions during training, denoted as ‘w/o’. The results in Fig. 4(b) demonstrate that the model’s performance decreases without rank rewards. Adding the amount $r$ of sample noise $\epsilon$ boosts the model’s performance. Specifically, $r=5$ strikes the optimal balance between personalization and semantic alignment.
5.5 Case Study
We present cases of personalized images generated by RAGAR and PMG across three datasets: POG, ML-latest, and Sticker. Fig. 5 shows eight historical images, a reference image, and generated images for context. The orange dashed boxes highlight the relevant items retrieved by the RAGAR retrieval module, arranged together for clearer visualization. In clothing scenario, RAGAR generates red-and-blue patterned shoes, effectively capturing the user’s preference for shoes while filtering out irrelevant items, such as patent leather shoes, thereby outperforming PMG in focusing on user preferences. In the movie poster scenario, RAGAR reflects preferences for multiple characters appearing together and avoids irrelevant anime aesthetics, which PMG fails to exclude. In the sticker scenario, RAGAR maintains semantic coherence between a cat and a guitar while aligning with the user’s preference for pink ears and a brown cat style. It successfully filters out realistic style preferences, demonstrating its ability to balance personalization and semantics. Additional cases are in the supplementary material.
Methods POG ML-latest
R@10 N@10 R@10 N@10
ORI 0.1765 0.1544 0.3229 0.2653
PMG 0.1804 0.1550 0.3213 0.2682
RAGAR 0.1927 0.1668 0.3299 0.2882
Table 4: Quantitative comparison of the recommendation performances.
5.6 Auxiliary Generation(RQ4)
Generating items beyond existing data can uncover user interests outside current data distribution [?]. Personalized images generated by RAGAR not only excel in display quality but also enhance recommendation tasks. We validate this using the multi-modal recommendation model MICRO[?]. We conduct experiments on two datasets with real interactions: POG and ML-latest. During training of rank model, reference images are replaced with those generated by RAGAR, PMG and original images ORI, respectively. The results in Tab. 4 reveal that generating images beyond original data distribution improves accuracy of the recommendation model. Both PMG and RAGAR outperform ORI across all metrics, demonstrating the benefit of generative methods. Notably, RAGAR shows a stronger capability in modeling user preferences compared to PMG, achieving significant improvements in metrics across both datasets, 6.8% (Recall@10) and 7.6% (NDCG@10) in POG, 2.67% (Recall@10) and 7.4% (NDCG@10) in ML-latest.
6 Conclusions and Future Work
In this paper, we proposed RAGAR for personalized image generation that balances personalization and semantics. With retrieval-augmented preferences, multi-modal alignment, and reflection-based optimization, RAGAR generates high-quality images that enhance both user experience and downstream recommendation performance. Extensive experiments demonstrate our method outperforms five baselines. For future work, we intend to unify preferences and noise in generation process to further enhance personalization.

### v2

@@
-RAGAR: Retrieval Augment Personalized Image Generation Guided by Recommendation
+RAGAR: Retrieval Augmented Personalized Image Generation Guided by Recommendation
 Abstract
@@
 Although effective, existing methods face two main issues.
-First, existing methods treat all items in the user historical sequence equally when extracting user preferences, overlooking the varying semantic similarities between historical items and the reference item.
-Disproportionately high weights for low-similarity items distort users’ visual preferences for the reference item.
-Second, existing methods heavily rely on consistency between generated and reference images to optimize the generation, which leads to underfitting user preferences and hinders personalization.
-To address these issues, we propose Retrieval Augment Personalized Image GenerAtion guided by Recommendation (RAGAR).
+First, existing methods treat all items in the user’s historical sequence equally when extracting user preferences, overlooking the varying semantic similarities between historical items and the reference item.
+Disproportionately high weights for low-similarity items distort user visual preferences for the reference item.
+Second, existing methods heavily rely on consistency between generated and reference images to optimize generation, which leads to underfitting user preferences and hinders personalization.
+To address these issues, we propose Retrieval Augmented Personalized Image GenerAtion guided by Recommendation (RAGAR).
 Our approach uses a retrieval mechanism to assign different weights to historical items according to their similarities to the reference item, thereby extracting more refined users’ visual preferences for the reference item.
@@
 Extensive experiments and human evaluations on three real-world datasets demonstrate that RAGAR achieves significant improvements in both personalization and semantic metrics compared to five baselines.
-1 Introduction
+Introduction
 Personalized image generation has been widely applied in scenarios such as advertising systems and chat software.
-It aims to render reference images into preferred ones based on user visual preferences.
-Existing image generation methods [???] often produce similar outputs for different users with similar input, failing to meet users’ need for personalization.
-As a personalized generation method, [?] extracts user preferences from user historical sequences by treating each item equally, ignoring the varying semantic similarities between the reference item and historical items, such as semantic differences in categories.
+It aims to render reference images into preferred ones based on the user’s visual preferences.
+Existing image generation methods (Nichol et al.
+2022; Rombach et al.
+2021; Gal et al.
+2023) often produce similar outputs for different users with similar input, failing to meet users’ needs for personalization.
+As a personalized generation method, PMG (Shen et al.
+2024) extracts user preferences from user historical sequences by treating each item equally, ignoring the varying semantic similarities between the reference item and historical items, such as semantic differences in categories.
 This disproportionately amplifies the influence of irrelevant items, causing the extracted user preference to deviate from the true visual preferences on the reference item.
@@
 Hence, we propose the assumption that filtering the items semantically related to the reference item enhances user preference modeling.
-Figure 1: A comparative example shows the effect of retrieval and recommendation by contrasting generated results with and without them.
+Figure 1: Effect of retrieval and recommendation on image quality.
+Lower rank and higher score of RAGAR-generated images suggest a strong positive correlation between retrieval/recommendation and personalized generation quality.
+Figure 2: Comparison of user history sequence construction strategies on POG.
+Top-$k$ retrieved items (Ret) consistently outperform lower-ranked (Exp-Ret) and random (Random) ones, validating that high-ranking relevant items better capture user preferences.
 Another challenge lies in defining and evaluating personalization quantitatively.
-Current methods rely on either human evaluation or large multi-modal models [??], both of which are resource-intensive.
-Recent work [?] evaluates and rectifies generated images heavily based on their pixel similarity to the reference image, resulting in limited personalization.
-Inspired by advances in recommendation systems, which excel in learning and assessing personalized preferences, we propose using multi-modal recommendation models to evaluate generated images, shown in Fig.
+Current methods rely on either human evaluation or large multi-modal models (Cho et al.
+2024; Wu et al.
+2024), both of which are resource-intensive.
+Recent work (Shen et al.
+2024) evaluates and rectifies generated images heavily based on their pixel similarity to the reference image, resulting in limited personalization.
+Inspired by advances in recommendation systems, which excel in learning and assessing personalized preferences, we propose using multi-modal recommendation models to evaluate images, shown in Fig.
 1(b).
 These models provide an efficient and effective way to measure personalization.
-Moreover, by combining models and metrics from the recommendation domain, we validate in Sec.
-2 that: 1) items selected through retrieval are more closely aligned with user preferences than randomly selected items; 2) retrieving items semantically related to the reference item enhances the learning of user preferences compared to irrelevant items.
-Hence, the paper proposed the Retrieval Augment Personalized Image GenerAtion guided by Recommendation (RAGAR), comprising three key modules: 1) Retrieval Module contains a correlation unit to calculate the semantic similarity between historical and reference items and a fusion unit to fuse the filtered visual features with dynamic weight.
-2) Generation module applies a balance calibrator to combine general user preferences derived from the fine-tuned LLM and the retrieval-augmented user preferences to produce images.
-3) Reflection module calculates the loss to balance semantics and personalization, guided by the discriminator.
-In summary, the key contributions are as follows:
-  • We are the first to emphasize the relationship between historical items and the reference item. The corresponding assumptions are supported by data analysis on real-world datasets in Sec. 2.
-  • We propose a novel model RAGAR to generate personalized images.
+Moreover, by combining models and metrics from the recommendation domain, we propose and validate two assumptions: 1) items selected through retrieval are more closely aligned with user preferences than randomly selected items; 2) retrieving items semantically related to the reference item enhances the learning of user preferences compared to irrelevant items.
+We compare three types of user historical sequences on POG (Chen et al.
+2019) dataset: Ret contains the top-$k$ items selected with the retrieval method described in Sec.
+Retrieval Module.
+Exp-Ret (Expanded Retrieval) contains items ranked between $(k,2k]$ by the retrieval method.
+Random contains $k$ randomly selected items.
+The results are shown in Fig.
+2.
+The Random sequence has the lowest values in metrics, demonstrating that randomly selected items are less effective for preference modeling.
+Our first assumption holds because both Ret and Exp-Ret achieve higher metrics, suggesting that items related to the reference items are crucial for modeling preferences.
+Moreover, both metrics of Ret are higher than those of Exp-Ret, which implies the second assumption holds.
+Therefore, both of our assumptions hold in practical applications, motivating us to retrieve higher-ranked items to enhance the preferences modeling.
+More results can be found in the Appendix.
+Hence, we propose RAGAR (Retrieval Augmented Personalized Image Generation guided by Recommendation), comprising three key modules.
+The Retrieval module identifies semantically relevant historical items via a correlation unit and aggregates their visual features through a fusion unit to construct a retrieval-augmented preference.
+The Generation module extracts item keywords from the user sequence and integrates both detailed and global preferences, calibrated by the balance calibrator, to generate personalized images.
+Finally, the Reflection module applies multiple assessment signals to guide the model toward achieving a better balance between personalization and semantic fidelity.
+The main contribution of the paper can be summarized as follows:
+  • To our knowledge, we are the first to emphasise and explore the intrinsic relationship between historical and reference items.
+We hypothesise this relationship is crucial in capturing preferences more accurately and guiding the personalized generation process.
+To support this hypothesis, we conduct a comprehensive data analysis on real-world datasets, which reveals significant impact of item-to-item relevance on personalization quality.
+  • We propose a novel personalized image generation framework, RAGAR, which is specifically designed to address the dual objectives of semantic alignment and personalization.
 We utilize the retrieval assumption realized by calculating the semantic similarity between items to enhance semantic consistency.
-Moreover, we design a discriminator to refine and assess the generated images, ensuring higher personalization quality.
-  • Experiments on three real-world datasets demonstrate that our proposed method can achieve significant improvements in both personalization and semantic alignment compared to five competing methods.
-2 Assumptions Validation
-In this section, we validate our assumptions through data analysis on three real-world datasets: POG, ML-latest, and Stickers.
-[^1] Dataset descriptions are provided in Sec.
-5.1.
-Details can be found in Table 1.
-To evaluate the effectiveness of the retrieval method, we use the multi-modal recommendation model [?] as the ranking model.
-We compare three types of user historical sequences: 1) ‘Ret’ contains the top-$k$ items selected using the retrieval method described in Sec.
-4.3.
-2) ‘Exp-Ret’ (Expanded Retrieval) contains the items ranked between $(k,2k]$ by the retrieval method.
-3) ‘Random’ contains $k$ randomly selected items.
-To assess how well these sequences capture user preferences, we use Recall@10 and NDCG@10 as evaluation metrics.
-Figure 2: Assumption validation results across three datasets.
-The results shown in Fig.
-2.
-‘Random’ has the lowest values in metrics, demonstrating that randomly selected items are less effective for preference modeling.
-Our first assumption holds because both ‘Ret’ and ‘Exp-Ret’ achieve higher metrics, suggesting that items related to the reference items are crucial for modeling user preferences.
-Moreover, both metrics of ‘Ret’ are higher than those of ‘Exp-Ret’, which implies that the second assumption holds.
-Therefore, both of our assumptions hold in practical applications, motivating us to retrieve higher ranked items to enhance the user preferences modeling.
-3 Related Work
-Figure 3: The overview framework of the proposed RAGAR model.
-Personalized Image Generation.
-Advances in diffusion models (DMs) [?] have enabled high-quality image generation [????].
-However, personalized image generation where user preferences guide the generation process, remains underexplored.
-Methods like [?] and [?] rely on natural language instructions, while [???] combine user interactions and instructions to create personalized advertisements or product images.
-Recent works bridge DMs and large language models (LLMs) [???], such as [?].
-Building on this foundation, [?] incorporates user historical sequences to extract personalized features.
-However, [?] relies solely on consistency loss for optimization, overfitting the reference image’s features while neglecting personalized preferences.
-End-to-end multi-modal large language models (MLLMs) [???] show promise in personalizing images from textual instructions often overlook fine-grained preferences.
-Moreover, widely used evaluation metrics like FID [?], SSIM [?], LPIPS [?], and CLIP score [?] do not align well with human judgment, making them inadequate for assessing personalization.
-Meanwhile, evaluating personalization manually or using MLLMs [??] is inherently resource-intensive.
-Recommendation Systems.
-The goal of recommendation systems is to deliver personalized content that meets user preferences.
-These systems often rely on retrieval [??] and recommendation [???] techniques.
-Under the multi-modal setting, approaches like [?] integrate textual and visual features for enhanced commonsense reasoning, while [?] extends [?] with image understanding adapters.
-Specifically, [?] captures user-item relationships across modalities, enabling precise preference modeling for personalized recommendations.
-4 Method
-In this section, we formulate our task, introduce key notation, and outlines the RAGAR mainstream.
-4.1 Task Formulation
-We introduce the notation and task formulation for personalized image generation.
-For the user $u$, let the historical sequence be denoted as $\mathbf{S}_{u}=\{I_{1},I_{2},\cdots,I_{N-1}\}$, and let $I_{N}$ represent the reference item.
-Personalized image generation aims to produce an image that aligns with the user’s personalized preferences and preserves semantic consistency with $I_{N}$.
-Under the multi-modal setting, each item $I_{j}$ consists of an image and a text description, denoted as $\{img_{j},txt_{j}\}$.
-4.2 RAGAR
+Additionally, we incorporate a discriminator to assess and enhance image quality, promoting better personalization.
+  • Experiments on three diverse and real-world datasets demonstrate that RAGAR outperforms five strong baseline methods across multiple evaluation metrics.
+The empirical results highlight the effectiveness of our retrieval-based design and the discriminator.
+Methods
+Task Formulation
+We formulate the task of personalized image generation in a multi-modal setting.
+Let $u$ denote a user, and let their historical interaction sequence be represented as: $\mathbf{S}_{u}=\{I_{1},I_{2},\dots,I_{N-1}\}$, where each item $I_{j}=\{\text{img}_{j},\text{txt}_{j}\}$ is a multi-modal pair consisting of an image $\text{img}_{j}$ and a corresponding textual description $\text{txt}_{j}$.
+The final item $I_{N}={\text{img}_{N},\text{txt}_{N}}$ serves as the reference item, which specifies the semantic content for generation.
+The goal is to generate a personalized image $\hat{\text{img}}_{u}$ that reflects the user’s historical preferences captured in $\mathbf{S}_{u}$ and aligns semantically with the reference item $I_{N}$.
+Formally, the task is to learn a generation function:
+$\hat{\text{img}}_{u}=\mathcal{G}(\mathbf{S}_{u},I_{N})$ (1)
+such that $\hat{\text{img}}_{u}$ maximizes personalization fidelity with respect to $\mathbf{S}_{u}$ and semantic consistency with respect to $I_{N}$.
+RAGAR
 Our proposed method, RAGAR, is illustrated in Fig.
 3.
-To reduce the impact of irrelevant items, we introduce the retrieval module that calculates semantic similarity scores between the reference item and historical items.
-Then, the module fuses the visual features of the items with the weight corresponding to the scores to produce retrieval-augmented preference features.
-Next, the generation module employs a fine-tuned LLM to derive general preference features from multi-modal historical sequences, corrected by the retrieval-augmented preference.
-A diffusion-based generator then generate images conditioned by corrected preference.
-Finally, we design a reflection module to evaluate the generated images, balancing personalization and semantics to ensure high-quality results.
-4.3 Retrieval Module
-The retrieval module aims to integrate visual features of historical items that are semantically relevant to the reference item, including the correlation unit and the fusion unit.
-Correlation Unit.
-The correlation unit transforms image information from historical items into textual features to compute semantic similarity scores with the reference item.
-Given the historical sequence $\mathbf{S}_{u}$ and reference item $I_{N}$ for the user $u$, we extract the textual description from the item images using a caption model (e.g., BLIP-2 [?]):
-$cap_{i}=\textbf{Caption}(img_{i}),i=0,1,\cdots,N$ (1)
-where $cap_{i}$ represents the caption of item $I_{i}$. Next, the textual description is processed into high-dimensional semantic features with a text encoder, such as CLIP [?]:
-$\mathbf{E}^{sem}_{i}=\mathbf{Encoder}(cap_{i})$ (2)
-where $\mathbf{E}^{sem}_{i}$ denotes the high-level semantic representation of each caption.
-Then, we calculate the semantic similarity between historical items and the reference item using cosine similarity:
-$s_{i}=cosine(\mathbf{E}^{sem}_{i},\mathbf{E}^{sem}_{N}),i=0,1,\cdots,N-1$ (3)
-where $s_{i}$ represents the similarity score for each historical item relative to the reference item.
-The similarity scores are sorted in descending order, and the top-$k$ items form the retrieval sequence $\mathbf{S}^{ret}_{u}$, capturing the most relevant items in the user’s history.
-Fusion Unit.
-The fusion unit integrates preference features from the retrieval sequence by computing a weighted sum of visual features based on their semantic similarity scores, emphasizing high-similarity items while filtering out low-similarity ones.
-For the retrieval sequence $\mathbf{S}^{ret}_{u}$, we extract the visual features using a visual encoder:
-$\mathbf{E}^{vis}_{i}=\mathbf{Encoder_{V}}(img_{i}),i=0,1,\cdots,N-1$ (4)
-where $\mathbf{E}^{vis}_{i}$ denotes the high-level visual representation of each image.
-Then, the retrieval-augmented preference feature is computed as a weighted sum of these visual features and their corresponding similarity scores:
-$\mathbf{P}^{ret}=\sum_{i=1}^{k}\mathbf{E}^{vis}_{i}\cdot s^{\prime}_{i}$ (5)
-where $\mathbf{P}^{ret}$ represents the retrieval-augmented preference feature for the user $u$, and $s_{i}^{\prime}$ denotes normalized similarity scores.
-4.4 Generation Module
-The generation module bridges general user preferences with retrieval-augmented preferences to produce personalized images.
-It operates in three stages: 1) extracting general preference features using an LLM, 2) correcting the general preference features with the balance calibrator, and 3) generating personalized images using a diffusion-based generator.
-Prompt Construction.
+To mitigate the influence of irrelevant historical items, the retrieval module computes semantic similarity scores between the reference item and historical items by correlation unit.
+The fusion unit then fuses the visual features of the most relevant items, weighted by similarity scores, to construct a retrieval-augmented preference.
+The generation module employs a LLM to extract and summarise item keywords from the sequence.
+The LLM and CLIP are then used to derive detailed and global user preferences, respectively, which are subsequently integrated via a balance calibrator.
+A diffusion-based generator then synthesizes personalized images based on the general preference.
+Finally, the reflection module evaluates the generated images and updates the model by balancing semantic consistency and personalization through multi-dimensional assessments, thereby ensuring high-quality results.
+Figure 3: Overview of the RAGAR framework.
+The framework consists of three stages.
+In the Retrieval stage, semantic and visual features are extracted to identify relevant historical items, whose visual features are fused into the retrieval-augmented preference.
+The Generation stage uses a LLM, CLIP and the balance calibrator to extract and integrate detailed and global preferences, enabling personalized image generation.
+In the Reflection stage, the generated images are evaluated through ranking, calibration, and semantic alignment, with policy gradients guiding the model to balance personalization and semantic fidelity.
+Retrieval Module
+Given the user’s interaction history $\mathbf{S}_{u}$ and a reference item $I_{N}$, we compute the semantic relevance of each historical item.
+First, we transform items’ image into the caption ${cap}_{i}$ using a pretrained image-to-text model (e.g., BLIP-2 (Li et al.
+2023)).
+Each caption is then encoded into a high-dimensional semantic embedding via a text encoder such as CLIP (Radford et al.
+2021):
+$\mathbf{E}^{sem}_{i}=\mathbf{Enc}^{sem}(cap_{i}),\quad i=0,1,\cdots,N$ (2)
+We compute the semantic similarity between each historical item and the reference item using cosine similarity:
+$s_{i}=\frac{\sum\mathbf{E}^{sem}_{i}\mathbf{E}^{sem}_{N}}{\sqrt{\sum{\mathbf{E}^{sem}_{i}}^{2}}\,\sqrt{\sum{\mathbf{E}^{sem}_{N}}^{2}}},\quad i=0,1,\cdots,N-1$ (3)
+The top-$k$ most similar items form the retrieval sequence $\mathbf{S}^{ret}_{u}$, capturing the most relevant preference signals from the user history.
+To form a retrieval-augmented preference, we extract visual features of the selected items using a visual encoder (e.g., ViT (Dosovitskiy et al.
+2021)):
+$\mathbf{E}^{vis}_{i}=\mathbf{Enc}^{vis}(img_{i}),\quad i=0,1,\cdots,N-1$ (4)
+These top-k features are then aggregated using the normalized similarity scores $s^{\prime}_{i}$ as weights.
+The resulting vector $\mathbf{P}^{ret}$ captures a preference representation that emphasizes semantically aligned visual cues while suppressing irrelevant information.
+$\mathbf{P}^{ret}=\sum_{i=1}^{k}\frac{\exp(s^{\prime}_{i})}{\sum_{j=1}^{k}\exp(s^{\prime}_{j})}\cdot\mathbf{E}^{vis}_{i}$ (5)
+Generation Module
+The generation module bridges global preference with detailed preference to produce personalized images.
 To capture user preferences from interactions, we transform interacted items into structured textual descriptions suitable for LLM analysis.
-Specifically, we adopt the approach from [?] that summarizes items with concise keywords to enhance interpretability and reduce extraneous information.
-We construct a prompt $p_{k}$ (detailed in supplementary material) populated with item captions and text to extract concise keywords with an LLM, denoted as $\mathbf{LLM}_{k}$:
-$w_{i}=\mathbf{LLM}_{k}(p_{k}(cap_{i},txt_{i})),i=0,1,\cdots,N-1$ (6)
-where $w_{i}$ is the keyword list for item $I_{i}$.
-We filter out low-frequency keywords and retain the top-$n$ keywords as $w^{\prime}$.
+Specifically, we summary items with concise keywords to enhance interpretability and reduce extraneous information.
+We construct a prompt $p_{k}$ (all prompts are detailed in Appendix) populated with item text to extract concise keywords with an LLM (e.g., Qwen2.5 (Bai et al.
+2023)), denoted as $\phi_{k}$:
+$w_{i}=\phi_{k}(p_{k}(cap_{i},txt_{i})),i=0,1,\cdots,N-1$ (6)
+With the keyword list $w_{i}$ for item $I_{i}$, we filter out low-frequency keywords and retain the top-$n$ keywords as $w^{\prime}$.
 Next, we construct a prompt $p_{g}$ to capture the general user preference.
-To enhance the expressivity of the LLM, denoted as $\mathbf{LLM}_{g}$, for image generation, we expand its vocabulary by incorporating $L$ additional special tokens ${\left[IMG\{i\}\right]}_{i=1}^{L}$.
-The tokens and keywords $w^{\prime}$ are then populated to $p_{g}$ (detailed in supplementary material) to extract preferences.
-General preference Extraction.
-Given the prompt $p_{g}$, the output of $\mathbf{LLM}_{g}$ can be divided into two embeddings:
-$[\mathbf{E}^{txt};\mathbf{E}^{img}]=\mathbf{LLM}_{g}(p_{g})$ (7)
-where $\mathbf{E}^{txt}$ represents text-related embedding and $\mathbf{E}^{img}\in\mathbb{R}^{L\times d}$ encodes image-related features.
-To bridge the gap between textual description and visual representation in LLMs, we adopt the Modal Mapper to align $E^{img}$ with the image space following [?], a 4-layer encoder-decoder transformer with trainable queries and two linear layers.
-Then we concatenate the text-related embedding and the image-related embedding and obtain the multi-modal feature.
-$\mathbf{E}^{m}=[\mathbf{E}^{txt};\mathbf{ModalMapper}(\mathbf{E}^{img})]$ (8)
-To maximize keyword utilization, we use the text encoder to generate keyword feature $\mathbf{E}^{key}$. We then employ a cross-attention layer to integrate $\mathbf{E}^{m}$ with $\mathbf{E}^{key}$:
-$\mathbf{P}^{gen}=[Attn(\mathbf{E}^{m},\mathbf{E}^{key},\mathbf{E}^{key});\mathbf{E}^{key}]$ (9)
-where $Attn$ denotes the cross attention layer to fuse features. In this way, we obtain the general preference feature $P^{gen}$.
-Balance Calibrator.
+To enhance the expressivity of the LLM, denoted as $\phi_{g}$, for image generation, we expand its vocabulary by incorporating $L$ additional special tokens ${[\mathbf{IMG}\{i\}]}_{i=1}^{L}$.
+The tokens and keywords $w^{\prime}$ are then populated to $p_{g}$ to extract preferences.
+Given the prompt $p_{g}$, we divide the output into text-related embedding $\mathbf{E}^{txt}$ and $\mathbf{[IMG]}$-related features $\mathbf{E}^{img}\in\mathbb{R}^{L\times d}$:
+$\mathbf{E}^{txt};\mathbf{E}^{img}]=\phi_{g}(p_{g})$ (7)
+To bridge the gap between textual description and visual representation in LLMs, we adopt the Modal Mapper, a 4-layer encoder-decoder transformer with trainable queries and two linear layers, to align $E^{img}$ with the image space following Gill (Koh, Fried, and Salakhutdinov 2023).
+Then we concatenate the text-related embedding and the image-related embedding and obtain the detailed multi-modal feature.
+$\mathbf{E}^{d}=[\mathbf{E}^{txt};\phi_{m}(\mathbf{E}^{img})]$ (8)
+To maximize keyword utilization and capture the global features, we use the text encoder to generate keyword feature $\mathbf{E}^{g}$.
+We employ a cross-attention layer and a residual connection to integrate $\mathbf{E}^{d}$ with $\mathbf{E}^{g}$.
+In this way, we obtain the general preference feature $P^{gen}$.
+The generator is then used to generate personalized image.
+$\mathbf{P}^{gen}=[\mathrm{softmax}(\frac{\mathbf{E}^{d}{\mathbf{E}^{g}}^{\top}}{{\sqrt{d_{\mathbf{E}^{g}}}}})\mathbf{E}^{g};\mathbf{E}^{g}]$ (9)
+Reflection Module
+As the diffusion-based generator, propagating gradients backwards from the generated images is challenging (Shen et al.
+2024).
+To address this issue, we design a multi-assessment reflection module based on policy gradient (Sutton et al.
+1999), ensuring both user alignment and semantic consistency.
+Recommendation Reflection
+To provide personalized feedback on generated images, we leverage a pre-trained multi-modal ranking model (e.g., MICRO (Zhang et al.
+2023)) to evaluate images, reducing the reliance on expensive manual labeling.
+More details about the ranking model (RM) can be found in Appendix.
+Given a set of images $\{v^{ref},v^{glob},v^{gen}\}$, where $v^{ref}$ represents the reference image, $v^{glob}$ represents the image generated with the global preference and $v^{gen}$ represents the image generated by RAGAR.
+By substituting visual features from this set into the sequence of original items, the RM assigns the scores $\rho\in\{\rho^{ref},\rho^{glob},\rho^{gen}\}$ and rankings $rk\in\{1,2,3\}$ derived from the score for each item:
+$\{\rho,rk\}=\textbf{RM}(S_{u},v)$ (10)
+Intuitively, images that better align with the user’s personalization preference are expected to achieve higher scores and lower ranks.
+To enable gradient propagation from image-level feedback, we adopt a policy gradient-inspired approach.
+The underlying formulation is based on a Markov Decision Process (MDP), defined as $\langle\mathcal{S},\mathcal{A},\mathcal{R},\mathcal{P},\gamma\rangle$ (state/action spaces, reward, transition probability, discount factor).
+In RAGAR, we simplify the multi-step decision process into a single-step generation by setting $\mathcal{P}=\gamma=1.0$, while retaining the core policy gradient framework.
+Model parameters are treated as states, image generation conditioned on user preference as actions, and rank-based rewards as action values.
+Although the states do not transition explicitly, parameter updates implicitly induce state changes.
+To support multi-action exploration, we sample $r$ perturbations $\epsilon_{t}$, $t\in[1,r]$, from a Gaussian distribution over user preferences and compute the final reward as the mean rank score of the generated images.
+Since policy gradients require maximizing reward while other losses are minimized, we incorporate the reward as a penalty term by negating it.
+Specifically, we employ a reward function to guide the model toward generating images that reflect personalization preferences.
+We then accumulate rewards for each noise $\epsilon_{t}$ and calculate the $\mathcal{L}_{rank}$.
+$\mathcal{R}_{t}=\sum_{c\in\{ref,glob\}}\max\left(0,\rho^{c}-\rho^{gen}_{t}\right)+\delta$ (11)
+$\mathcal{L}_{rank}=-\frac{1}{r}\sum_{t=1}^{r}\log_{\pi_{\theta}}(\mathbf{P}^{gen}+\epsilon_{t})\cdot\mathcal{R}_{t}$ (12)
+To encourage the model to generate higher-quality images we add a margin $\delta$.
+And $\log_{\pi_{\theta}}$ computes the probability density of the generated embedding under the Gaussian distribution.
+Calibrator Reflection
 As mentioned in Sec.
-4.3, the retrieval-augmented preference focuses primarily on item features associated with the reference item.
-To narrow the gap between the retrieval-augmented preference feature $\mathbf{P}^{ret}$ and the global preference feature $\mathbf{P}^{gen}$, we minimize the calibrator loss between them:
+Retrieval Module, the retrieval-augmented preference focuses primarily on item features associated with the reference item.
+To narrow the gap between the general preference feature $\mathbf{P}^{gen}$ and the retrieval-augmented preference feature $\mathbf{P}^{ret}$, we minimize the calibrator loss between them:
 $\mathcal{L}_{\mathrm{cal}}=\|\mathbf{P}^{gen}-\mathbf{P}^{ret}\|_{2}^{2}$ (13)
-This adjustment ensures that $P^{gen}$ reflects both the general and retrieval-augmented preferences.
-Image Generation.
-With the corrected preference, we utilize the diffusion-based generator to generate images.
-During training, we sample $r$ random noises $\{\epsilon_{t}\}_{t=1}^{r}\sim\mathcal{N}(\mu,\sigma^{2}I)$ and apply reparameterization trick to enable gradient backpropagation through the sampling process:
-$\mathbf{z_{t}}=\mu+\sigma\odot\epsilon_{t}$ (11)
-The generator then produces $r$ images using the policy gradient update strategy discussed in Sec.
-4.5.
-Additionally, the keyword feature $\mathbf{E}^{key}$ is used to generate a reference-like image, which serves as input to the ranking model for comparison with the personalized images during the reward computation.
-$\begin{split}v^{key}&=Unet(\mathbf{E}^{key})\\ v^{gen}_{t}&=Unet(\mathbf{P}^{gen}+\mathbf{z_{t}}),t=0,1,...,r\end{split}$ (12)
-where $v^{key}$ represents the image generated using the keyword features, and $v^{gen}$ represents the image generated using the general preference features.
-4.5 Reflection Module
-As the diffusion-based generator, propagating gradients backwards from the generated images is challenging [?].
-To address this issue, we design a two-part reflection module that balances personalization and semantics in the generated images, ensuring both user alignment and semantic consistency.
-Personalization Reflection
-To provide personalized feedback on generated images, we leverage a pre-trained multi-modal ranking model (RM) to evaluate images, reducing the reliance on expensive manual labeling.
-Given a set of images $\{v^{ref},v^{key},v^{gen}_{t}\}$, where $v^{ref}$ represents the reference image.
-By substituting visual features from this set into the sequence of original items, RM assigns the scores and rankings for each item:
-$\{\rho,rk\}=\textbf{RM}(S_{u},v)$ (13)
-where $\rho\in\{\rho^{ref},\rho^{key},\rho^{gen}_{t}\}$ is the rank score, and $rk\in\{1,2,3\}$ denotes the rank derived from the score.
-Intuitively, images that align better with the user’s personalization preference are expected to achieve higher scores and lower ranks.
-We then employ a reward function to guide the model toward generating images that reflect personalization preferences.
-Inspired by the policy gradient method [?], we accumulate rewards for each noise $\epsilon_{t}$:
-$\mathcal{R}_{t}=\sum_{c\in\{ref,key\}}\max\left(0,\rho^{c}-\rho^{gen}_{t}\right)+\delta$ (14)
-$\mathcal{L}_{rank}=-\frac{1}{r}\sum_{t=1}^{r}\log_{\pi_{\theta}}(\mathbf{P}^{gen}+\epsilon_{t})\cdot\mathcal{R}_{t}$ (15)
-where $\delta$ is a margin that encourages the model to generate higher-quality images, $\log_{\pi_{\theta}}$ computes probability density of the generated embedding under the Gaussian distribution.
+This adjustment ensures that $P^{gen}$ reflects both the general and retrieval-augmented preferences, and guide the model to learn from retrieval.
 Semantic Reflection
-To enhance the semantic consistency, we minimize the distance between the mapper’s output in the generation module, $\mathbf{E}^{m}$, and the semantic features of the reference image, $\mathbf{E}^{sem}_{N}$ using the semantic loss:
-$\mathcal{L}_{\mathrm{sem}}=\|\mathbf{E}^{m}-\mathbf{E}^{sem}_{N}\|_{2}^{2}$ (16)
-Combined with the calibrator loss, generated images align both semantically and visually with the reference.
-Joint Reflection
-We employ LoRA [?] to update a limited set of parameters in LLM, the Modal Mapper and the attention fusion components. The overall loss is defined as:
-$\mathcal{L}=\alpha\cdot\mathcal{L}_{cal}+\beta\cdot\mathcal{L}_{\mathrm{rank}}+\gamma\cdot\mathcal{L}_{\mathrm{sem}}$ (17)
-where $\alpha$, $\beta$ and $\gamma$ are adjustable weighting hyper-parameters.
-Features POG ML-latest SER30K
-#Users 1000 4689 2230
-#Items 19242 9742 30739
-Table 1: The characteristics of experimental datasets.
-5 Experimentation
-We evaluate RAGAR on three different scenarios: commodity, movie poster and sticker. Our experiments aim to answer the following research questions:
+To enhance the semantic consistency, we minimize the distance between the modal mapper’s output in the generation module, $\mathbf{E}^{d}$, and the semantic features of the reference image, $\mathbf{E}^{sem}_{N}$, using the semantic loss.
+Combined with the semantic loss, generated images align more semantically with the reference.
+$\mathcal{L}_{\mathrm{sem}}=\|\mathbf{E}^{d}-\mathbf{E}^{sem}_{N}\|_{2}^{2}$ (14)
+Reflection Loss
+We update the Balance Calibrator in the generation module. The overall loss is defined with three adjustable weighting hyper-parameters $\alpha$, $\beta$ and $\gamma$:
+$\mathcal{L}=\alpha\cdot\mathcal{L}_{rank}+\beta\cdot\mathcal{L}_{\mathrm{cal}}+\gamma\cdot\mathcal{L}_{\mathrm{sem}}$ (15)
+Experiment
+We evaluate RAGAR on three different scenarios: commodity, movie poster, and sticker. Our experiments aim to answer the following research questions:
   • RQ 1: How does RAGAR perform compared to other generative methods in quantitative evaluation metrics?
-  • RQ 2: Dose RAGAR outperform baseline methods in human evaluation?
-  • RQ 3: What is the impact of each module of RAGAR on performance? Specifically, how do the retrieval module and the reflection module contribute in the generation?
+  • RQ 2: Does RAGAR outperform baseline methods in human evaluation?
+  • RQ 3: What is the impact of each module of RAGAR on performance? Specifically, how do the retrieval module and the reflection module contribute to the generation?
   • RQ 4: Can personalized images generated by RAGAR improve the performance in recommendation systems?
-5.1 Experimental Settings
-Datasets.
-We utilize three real-world datasets to generate personalized images across different scenarios: POG [^2] https://github.com/wenyuer/POG.,ML-latest [^3] https://grouplens.org/datasets/movielens.,SER30K [^4] https://github.com/nku-shengzheliu/SER30K..
-POG is a multi-modal dataset of fashion clothing with user interaction history.
-Due to the large size, we randomly select a subset of users and interactions.
-ML-latest is a benchmark movie dataset with user ratings.
-We collect corresponding posters from IMDB [^5] https://www.imdb.com.
-and split user historical sequences to ensure each contains 20 interactions.
-SER30K is a large-scale sticker dataset where each sticker is categorized by theme and annotated with an associated emotion label.
-We merge two random themes into user historical sequences and randomly select a reference sticker from other themes.
-Dataset statistics are summarized in Table 1.
 Datasets Methods Personalization Semantic Alignment
@@
 Table 1: Quantitative performance comparison on three datasets in terms of personalization and semantic alignment. The best performance is bold while the second-best is underlined.
+Experimental Settings
+Features POG ML-latest SER30K
+#Users 1000 4689 2230
+#Items 19242 9742 30739
+Table 2: Characteristics of experimental datasets.
+Datasets.
+We utilize three real-world datasets to generate personalized images across different scenarios: POG, ML-latest ,SER30K.
+POG is a multi-modal dataset of fashion clothing with user interaction history.
+Due to the large size, we randomly select a subset of users and interactions.
+ML-latest is a benchmark movie dataset with user ratings.
+We collect corresponding posters from IMDB and split user historical sequences to ensure each contains 20 interactions.
+SER30K is a large-scale sticker dataset where each sticker is categorized by theme and annotated with an associated emotion label.
+We merge two random themes into user historical sequences and randomly select a reference sticker from other themes.
+Dataset statistics are summarized in Tab.
+2.
 Comparison Methods.
-We compare RAGAR with five generative baselines, including three DM-based models Glide [?], SD [?] and TI [?], and two LLM-based models LaVIT [?] and PMG [?].
+We compare RAGAR with five generative baselines, including three DM-based models Glide (Nichol et al.
+2022), SD (Rombach et al.
+2021), and TI (Gal et al.
+2023), and two LLM-based models LaVIT (Jin et al.
+2024) and PMG (Shen et al.
+2024).
 Evaluation Metrics.
-We compare RAGAR with baselines through the following quantitative metrics: To evaluate personalization, we calculate the CLIP Personalization Score (CPS) and the CLIP Personalization Image Score (CPIS), which measure the similarity between the generated images and the text description or images representing user preferences, respectively.
-In addition, we calculate the LPIPS [?] and SSIM [?] to quantify the perceptual similarity.
+We compare RAGAR with baselines through the following quantitative metrics: To evaluate semantic alignment, we calculate CLIP Score (CS), CLIP Image Score (CIS), LPIPS (Zhang et al.
+2018) and SSIM (Wang et al.
+2004) for reference images.
+To evaluate personalization, we calculate the CLIP Personalization Score (CPS) and the CLIP Personalization Image Score (CPIS), which measure the similarity between the generated images and the text description or images representing user preferences, respectively.
+In addition, we calculate the LPIPS (Zhang et al.
+2018) and SSIM (Wang et al.
+2004) to quantify the perceptual similarity.
 Furthermore, we compute the rank change $\mathbf{\Delta R}$ between the original and generated images.
@@
 A larger $\mathbf{\Delta R}$ indicates greater personalization in the image.
-$\mathbf{\Delta R}=\frac{rk_{ori}-rk_{gen}}{1+rk_{ori}}$
-where $rk$ denoted the rank determined by the ranking model. To evaluate semantic alignment, we calculate CLIP Score (CS), CLIP Image Score (CIS), LPIPS [?] and SSIM [?] for reference images.
+$\mathbf{\Delta R}=\frac{rk_{ori}-rk_{gen}}{1+rk_{ori}}$ (16)
 Parameter Settings.
@@
 The number of noise is set at 3.
+The hyperparameters for joint reflection are $\alpha=0.2,\beta=0.5,\gamma=0.3$.
+More training details can be found in the Appendix.
 All experiments are conducted on a single NVIDIA-A100 GPU.
-5.2 Performance Comparison (RQ1)
-The comparison between RAGAR and baselines is shown in Tab.
-2.
-We can draw the following conclusions.
-First, traditional DM-based methods perform poorly on personalization metrics.
-Specifically, Glide and SD are the worst performers across five personalization metrics in all datasets.
-This is because both methods rely on the CLIP model to generate images, which struggles to capture deeper associations between texts.
-TI introduces a word embedding approach based on SD to capture user preferences.
-The stylized word embedding significantly improves personalization performance across all datasets (e.g., for $\Delta R$: -25.64 vs.
+Performance Comparison (RQ1)
+Tab.
+1 presents the comparison between RAGAR and baseline methods, from which several key observations can be made.
+First, traditional DM-based methods show poor performance on personalization metrics.
+Glide and SD rank lowest across all five personalization metrics on every dataset, largely because both rely on CLIP for image generation, which struggles to capture deeper semantic associations.
+TI, which augments SD with stylized word embeddings to model user preferences, significantly improves personalization (e.g., $\Delta R$: -25.64 vs.
 -54.02, -2.58 vs.
@@
 0.86).
-In terms of semantic alignment, all three methods perform well, and Glide achieves optimal results on SSIM across all datasets.
-This indicates that traditional methods tend to focus on reconstructing reference images.
-Second, LLM-based methods demonstrate better performance, benefiting from their ability to understand and analyze text comprehensively.
-LaVIT, a multi-modal vision-language transformer, is designed to align textual and visual features effectively.
-Although it outperforms DM-based methods, it still falls short in personalization ($\Delta R$: -51.09 in the POG dataset).
-PMG, an LLM-based personalized image generation method, leverages LLMs to extract user preferences.
-This approach significantly enhances personalization performance, achieving optimal results compared to DM-based baselines.
-However, it heavily relies on the consistency with reference images, limited the personalization performance.
-Third, our proposed method, RAGAR, achieves state-of-the-art (SOTA) performance due to the retrieval for sequence and the rank model for training across the three datasets.
-The performance boost is more pronounced on the ML dataset compared to POG and Sticker, likely because the amount of data in the front is significantly larger than that in the back.
-Compared to PMG, the second-best model, our method enhances consistency by semantically retrieving sequences and introducing evaluation metrics that better align with human judgment, resulting in stronger personalization performance.
+In terms of semantic alignment, all three methods perform well, with Glide achieving the highest SSIM, indicating a tendency to reconstruct reference images.
+Second, LLM-based methods perform better overall, owing to their stronger textual understanding.
+LaVIT, a multi-modal transformer, better aligns visual and textual features but still underperforms in personalization (e.g., $\Delta R$: -51.09 on POG).
+PMG, which extracts user preferences via LLMs, achieves better personalization than DM-based baselines but remains constrained by its dependence on reference image consistency.
+Finally, RAGAR achieves state-of-the-art performance across all datasets, benefiting from sequence retrieval and a ranking-based training strategy.
+The gain is especially notable on the ML dataset, likely due to its larger volume of front-end user data.
+Compared to PMG, RAGAR further improves consistency through semantic retrieval and the use of evaluation metrics more aligned with human judgment, leading to superior personalization.
 Methods POG ML-latest Sticker
@@
 ORI 2.26 / 2.16 / 2.18 /
-PMG 2.08 1.62 2.1 1.94 2.24 1.88
+PMG 2.08 1.62 2.10 1.94 2.24 1.88
 RAGAR 1.66 1.46 1.74 1.66 1.58 1.12
-Table 3: Human evaluation comparison on both personalization and semantic alignment.
-5.3 Human Evaluation (RQ2)
-To assess the personalization and semantics of RAGAR in real scenarios, we conduct human evaluation to compare it with PMG and original images.
-We invited 50 volunteers and set up two types of sorting tasks, with 50 cases for each dataset, covering personalization and semantics.
-For personalization, given 5 historical images along with the generated images, participants are asked to sort images in descending order of likelihood that the user would click them, based on historical click data.
-For semantics, given the reference image and generated images, participants are asked to sort them in descending order of how well they convey the same meaning as the reference image.
+Table 3: Human evaluation of RAGAR, PMG and origin images.
+RAGAR consistently outperforms both ORI and PMG in all metrics, achieving up to 29.46% improvement in personalization and 40.43% in semantic alignment.
+Human Evaluation (RQ2)
+To evaluate RAGAR’s effectiveness in personalization and semantic alignment, we conducted a human study comparing it with PMG and original images.
+Fifty participants completed two ranking tasks (50 cases per dataset): one on personalization—ranking generated images by their likelihood of being clicked based on 5 historical items—and one on semantics—ranking images by how well they convey the meaning of a reference image.
 As shown in Tab.
-3, RAGAR outperforms the baselines on both metrics.
-This indicates RAGAR better reflects user preference while preserving the semantics.
-Figure 4: Ablation study of the retrieval module and the rank rewards.
-Figure 5: Examples of image generation across three datasets.
-5.4 Ablation Study (RQ3)
-We study the effectiveness of key components of RAGAR in the POG dataset, including retrieval and reflection module.
+3, RAGAR consistently outperforms baselines on both tasks, demonstrating its superior ability to capture user preferences while preserving semantic content.
+Ablation Study (RQ3)
+We study the effectiveness of key components of RAGAR in the POG dataset, including the retrieval and reflection module.
+Experiments on more datasets can be found in the Appendix.
 Effect of retrieval module.
-We assess the impact of retrieval module on improving semantic consistency and capturing user preferences by excluding it during training.
+We assess the impact of the retrieval module on improving semantic consistency and capturing user preferences by excluding it during training.
 Fig.
 4(a) shows that: 1) noisy items in the historical sequence diminish the model’s ability to capture user preferences and interfere with semantics; 2) the retrieval module selects items relevant to user preferences, thereby improving the performance.
-Furthermore, we investigate the effect of varying retrieval number $k$ during training.
-The results reveal that $k=5$ yields the best results when historical sequence length is 20.
+Furthermore, we investigate the effect of varying the retrieval number $k$ during training.
+The results reveal that $k=5$ yields the best results when the historical sequence length is 20.
 At $k=0$ (no retrieval), performance is poor, while increasing $k$ initially improves performance but declines beyond $k=5$ due to reintroduced noise and redundancy.
 Effect of rank rewards.
-We investigate the role of rank rewards by excluding it while retaining the other loss functions during training, denoted as ‘w/o’.
+We investigate the role of rank rewards by excluding it while retaining the other loss functions during training, denoted as w/o.
 The results in Fig.
@@
 Specifically, $r=5$ strikes the optimal balance between personalization and semantic alignment.
-5.5 Case Study
-We present cases of personalized images generated by RAGAR and PMG across three datasets: POG, ML-latest, and Sticker.
-Fig.
-5 shows eight historical images, a reference image, and generated images for context.
-The orange dashed boxes highlight the relevant items retrieved by the RAGAR retrieval module, arranged together for clearer visualization.
-In clothing scenario, RAGAR generates red-and-blue patterned shoes, effectively capturing the user’s preference for shoes while filtering out irrelevant items, such as patent leather shoes, thereby outperforming PMG in focusing on user preferences.
-In the movie poster scenario, RAGAR reflects preferences for multiple characters appearing together and avoids irrelevant anime aesthetics, which PMG fails to exclude.
-In the sticker scenario, RAGAR maintains semantic coherence between a cat and a guitar while aligning with the user’s preference for pink ears and a brown cat style.
-It successfully filters out realistic style preferences, demonstrating its ability to balance personalization and semantics.
-Additional cases are in the supplementary material.
+Figure 4: The impact of the retrieval module and the rank reward mechanism.
+Performance improves as the number of retrieval items and noise samples increases, peaking at moderate values (around 5 for retrieval and noise), after which excessive inputs introduce noise and degrade results.
 Methods POG ML-latest
@@
 RAGAR 0.1927 0.1668 0.3299 0.2882
-Table 4: Quantitative comparison of the recommendation performances.
-5.6 Auxiliary Generation(RQ4)
-Generating items beyond existing data can uncover user interests outside current data distribution [?].
-Personalized images generated by RAGAR not only excel in display quality but also enhance recommendation tasks.
-We validate this using the multi-modal recommendation model MICRO[?].
-We conduct experiments on two datasets with real interactions: POG and ML-latest.
-During training of rank model, reference images are replaced with those generated by RAGAR, PMG and original images ORI, respectively.
-The results in Tab.
-4 reveal that generating images beyond original data distribution improves accuracy of the recommendation model.
-Both PMG and RAGAR outperform ORI across all metrics, demonstrating the benefit of generative methods.
-Notably, RAGAR shows a stronger capability in modeling user preferences compared to PMG, achieving significant improvements in metrics across both datasets, 6.8% (Recall@10) and 7.6% (NDCG@10) in POG, 2.67% (Recall@10) and 7.4% (NDCG@10) in ML-latest.
-6 Conclusions and Future Work
-In this paper, we proposed RAGAR for personalized image generation that balances personalization and semantics.
-With retrieval-augmented preferences, multi-modal alignment, and reflection-based optimization, RAGAR generates high-quality images that enhance both user experience and downstream recommendation performance.
-Extensive experiments demonstrate our method outperforms five baselines.
-For future work, we intend to unify preferences and noise in generation process to further enhance personalization.
+Table 4: Application of personalized generation in recommendation domain. Compared to PMG and original images(ORI), RAGAR leads to an average improvement of approximately 7% in recommendation metrics.
+Figure 5: Qualitative comparison of different methods.
+Three representative cases highlight how RAGAR incorporates retrieved features into the generation.
+Left: user interaction history (dashed boxes indicate retrieved items).
+Right: images generated by RAGAR and PMG.
+Auxiliary Generation (RQ4)
+Generating items beyond the original data distribution can reveal user interests not captured by existing data (Li et al.
+2024).
+RAGAR-generated personalized images not only offer high visual quality but also improve recommendation performance.
+We validate this using the multi-modal recommendation model MICRO (Zhang et al.
+2023) on two real-world datasets: POG and ML-latest.
+In our experiments, we replace reference images with those generated by RAGAR, PMG, and original images (ORI) during rank model training.
+As shown in Tab.
+4, both PMG and RAGAR outperform ORI across all metrics, highlighting the advantage of generative approaches.
+RAGAR achieves the best results, improving Recall@10 and NDCG@10 by 6.8% and 7.6% on POG, and by 2.67% and 7.4% on ML-latest, demonstrating its superior ability to model user preferences.
+Case Study
+We showcase personalized image generation results from RAGAR and PMG on the POG, ML-latest, and Sticker datasets (Fig.
+5).
+Each example includes eight historical items, a reference image, and generated outputs, with orange dashed boxes marking items retrieved by RAGAR.
+In the clothing case, RAGAR produces red-and-blue shoes that match user preferences while filtering out irrelevant styles like patent leather, outperforming PMG.
+In the movie poster case, RAGAR reflects the user’s preference for multiple characters and avoids unrelated anime elements, which PMG fails to do.
+In the sticker case, RAGAR maintains semantic consistency (e.g., cat and guitar) and style (e.g., pink ears, brown fur), while excluding unwanted realistic styles.
+More examples are in the Appendix.
+Related Works
+Diffusion models (DMs)
+(Rombach et al.
+2021) have enabled high-quality image generation (Nichol et al.
+2022; Xu et al.
+2024b; Wei et al.
+2024; Zeng et al.
+2024; Wang et al.
+2024; Zhang et al.
+2025; He et al.
+2025; Lu et al.
+2025; Bi et al.
+2024; Zhang et al.
+2024) and video generation (Bi et al.
+2025; Xu et al.
+2025; Wang et al.
+2025a; Cao et al.
+2025; Feng et al.
+2024; Shao et al.
+2025), but personalization remains underexplored.
+Early methods (Gal et al.
+2023; Ruiz et al.
+2023) rely on text prompts, while others (Xu et al.
+2024a; Yang et al.
+2024; Shilova et al.
+2023) combine user interactions and prompts to generate personalized content.
+Some approaches connect DMs with large language models (LLMs) (OpenAI et al.
+2024; Touvron et al.
+2023; Bai et al.
+2023), such as (Koh, Fried, and Salakhutdinov 2023), and (Shen et al.
+2024).
+However, (Shen et al.
+2024) depends solely on consistency loss, often overfitting the reference image and ignoring user preferences.
+Multi-modal LLMs (Liu et al.
+2023; Anil et al.
+2023; Jin et al.
+2024) generate personalized images from text but lack fine-grained preference modeling.
+Recommendation systems
+aim to deliver content aligned with user preferences, using retrieval (Borgeaud et al.
+2022; Cui et al.
+2024; Wang et al.
+2025b) and ranking methods (Zhao et al.
+2024; Geng et al.
+2023; Zhang et al.
+2023).
+Multi-modal approaches like (Cui et al.
+2024) fuse text and image features for better reasoning, while (Geng et al.
+2023) adds visual understanding.
+(Zhang et al.
+2023) further models cross-modal user-item relationships.
+More detailed related work provide in the Appendix.
+Conclusion
+In the paper, we propose and validate two key assumptions in Sec.
+Introduction.
+Based on these insights, we introduce RAGAR, a novel framework that enhances personalization via retrieval-augmented preference and recommendation while preserving semantic consistency.
+We propose the Balance Calibrator to combine global and detailed preferences for guiding personalized generation.
+To address gradient propagation at the image level, we integrate recommendation signals and policy gradients into training.
+In future work, we aim to inject user preferences directly into the diffusion process to better align generation with user intent.